OpenStudy (anonymous):

Use the fact that the cosine is even and the sine is odd to determine if the following are even, odd, or neither.

OpenStudy (anonymous):

\[\cot x\]

OpenStudy (anonymous):

@ParthKohli @satellite73

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (aravindg):

cot x=cos x/sin x even/odd so odd

OpenStudy (anonymous):

How do I show that?

OpenStudy (aravindg):

Take an example of x^2 (even) and and x(odd). x^2/x=x which is odd

OpenStudy (aravindg):

Does that make sense?

OpenStudy (anonymous):

ohh okay so with this... does this make sense \[\cot(- x)= \frac{ \cos (-x) }{ \sin (-x) }\]

OpenStudy (aravindg):

Yeah that is the actual method.

OpenStudy (anonymous):

cot is actually odd... \(\cot x=\Large{\frac{\cos x}{\sin x}}\) so \(\cot (-x)=\Large{\frac{\cos (-x)}{\sin (-x)}}=-\Large{\frac{\cos (x)}{\sin (x)}}=-\cot x\)

OpenStudy (aravindg):

you get cot(-x)=-cot x

OpenStudy (anonymous):

WOW! you guys are so helpful! Thanks @AravindG and @pgpilot326 !! Who should I give my medal too!

OpenStudy (aravindg):

yw :)

OpenStudy (anonymous):

no worries... good luck!

OpenStudy (anonymous):

I'll give one, too. I think you can multiply and divide such that even functions are +1 and odd are -1 and the results will be + or - as appropriate. cot = cos/sin = (1)/(-1) = -1 thus odd.

OpenStudy (anonymous):

thanks @douglaswinslowcooper ! that's another good way.